Two railroad cars, each of mass 58000 kg kg and traveling 85 km/h km/h in opposite directions, collide head-on and come to rest.
Answer #1

m = 58000kg

v = 85 km/h = 23.6 m/s

the energy comes from the conversion of kinetic energy, so

KE=1/2 mv² = 1/2 (5.8x104 kg)(23.6 m/s)² = 16151840 J

the combination produces double kinetic energy

Therefore total kinetic energy is = 2 x16151840 J = 32303680J = 3.23 x 107J